OpenStudy (anonymous):

I am messing up this problem: Multiply and simplify: √7x(√x - 7√7). 1. √7+(√x) + √7x(-7√7) 2. √7x2 + √x(-7) 3. ?

Nnesha (nnesha):

^7x ?? what's this sign "^" for ?

OpenStudy (anonymous):

square root

Nnesha (nnesha):

lel okay so \[\sqrt{7x} (\sqrt{x - 7^7})\] like this ?? :o can you use drawing tool plz :-)

Nnesha (nnesha):

http://prntscr.com/6y63to

OpenStudy (anonymous):

√7x(√x-7√7) this is the original problem

OpenStudy (wmj259):

\[\sqrt{7x}(\sqrt{x}-7\sqrt{7})\] Is that right?

OpenStudy (anonymous):

yes

OpenStudy (shiraz14):

@Rose2812: If the question was as indicated by wmj259, i.e.: √(7x)(√x - 7√7) then the answer would be: (√7)x - 49√x

OpenStudy (anonymous):

Thank you for your answer but I am also looking for the steps to get there.

Nnesha (nnesha):

okay so question is \[\huge\rm \sqrt{7x} (\sqrt{x} - 7 \sqrt{7})\] distribute parentheses by sqrt{7x}

Nnesha (nnesha):

\[\huge\rm \sqrt{7x} \times \sqrt{x} - \sqrt{7x}\color{red}{ 7}\sqrt{7}\] first term of 2nd step is right for 2nd one multiply 7x by 7(which is under the square root ) red 7 would stay outside the root

Nnesha (nnesha):

let me know if you have aany question about it :-)

OpenStudy (anonymous):

Thank you!

Nnesha (nnesha):

did you get the answer ?

OpenStudy (anonymous):

Yes, now I just have to figure out how to input exponential functions on a graphing calculator!

Nnesha (nnesha):

ohh cool! :-)